### v1

# TFPI: Thinking-Free Policy Initialization Makes Distilled Reasoning Models More Effective and Efficient Reasoners

Abstract
Reinforcement Learning with Verifiable Reward (RLVR) effectively solves complex tasks but demands extremely long context lengths during training, leading to substantial computational costs. While multi-stage training can partially mitigate this, starting with overly short contexts often causes irreversible performance degradation, ultimately failing to reduce overall training compute significantly. In this paper, we introduce Thinking-Free Policy Initialization (TFPI), a simple yet effective adaptation to RLVR that bridges long Chain-of-Thought (CoT) distillation and standard RLVR. TFPI employs a simple ThinkingFree operation, explicitly discarding the thinking content via a direct </think> append, to reduce token usage during inference. Training with ThinkingFree-adapted inputs improves performance and lowers token consumption, even in the original slow-thinking mode. Extensive experiments across various benchmarks have shown that TFPI accelerates RL convergence, achieves a higher performance ceiling, and yields more token-efficient reasoning models without specialized rewards or complex training designs. With TFPI only, we train a 4B model to reach 89.0% accuracy on AIME24 and 65.5% on LiveCodeBench using less than 4K H20 hours.
footnotetext: $\dagger$ Corresponding Authors.
[Uncaptioned image] Github Repo [GitHub Page]
[Uncaptioned image] Models [Huggingface Models]
1 Introduction
Reasoning is a fundamental aspect of human cognition, and equipping artificial intelligence (AI) with strong reasoning capabilities is critical for its deployment and applications (Morris et al., 2023; Huang et al., 2024; Xu et al., 2025c). Progress in pretraining (Shao et al., 2024; Yang et al., 2024; Chen et al., 2025b), supervised fine-tuning (SFT) (Yu et al., 2023; Xu et al., 2024; Tong et al., 2024; Ye et al., 2025; Muennighoff et al., 2025), rigorous evaluation (Rein et al., 2024; Phan et al., 2025; Xu et al., 2025b), reinforcement learning (RL) (Jaech et al., 2024; Guo et al., 2025) has significantly enhanced the reasoning abilities of large language models (LLMs). Among these, RL with verifiable rewards (RLVR) stands out as an effective approach that enables large language models (LLMs) to generate long Chains-of-Thought (CoT) (Wei et al., 2022) spontaneously, and empowers them with unprecedented performance on challenging reasoning tasks. Thus, these RLVR-trained LLMs are termed as long-CoT LLMs, slow-thinking LLMs, or large reasoning models (LRMs).
Compared with initializing from a base LLM, starting from an SFT-distilled LRM typically yields better results and accelerates convergence in RLVR (Guo et al., 2025; An et al., 2025). However, SFT-distilled LRMs often produce excessively long responses during the rollout stage of RLVR, which necessitates a large training context length for RLVR. Using such large training contexts also incurs substantial computational costs. A common mitigation strategy is multistage RLVR, which begins with a relatively “short” context and gradually increases the training length (Luo et al., 2025b; An et al., 2025). Nonetheless, Zeng et al. (2025a) argue that multistage training might cause irreversible performance degradation. Moreover, even multistage training demands significant computational resources. For instance, training a 4B model while progressively increasing the maximum context length from 40K to 48K to 52K tokens requires approximately 8K H800 GPU hours (An et al., 2025). These limitations underscore the need for more efficient RLVR training methods.
Figure 1: Our proposed TFPI accelerates the convergence of RLVR to a higher performance ceiling (left) and yields more token-efficient reasoning models (right). Left: avg@32 versus training compute, measured in H20 hours. “Direct RL” refers to directly training Qwen3-4B with a 32K context window using DAPO, while “TFPI + RL” denotes running 32K-context DAPO after initialization with our 3-stage TFPI. The x-axis for TFPI uses a linear scale during the TFPI phase, followed by a logarithmic scale, with the transition indicated by a black vertical line. Right: Average accuracy on 4 reasoning datasets (AIME24/25, Beyond AIME, and GPQA) versus average output tokens. Points in the upper-left region indicate better performance. Baseline names and their corresponding numbers are listed in Table 3. Red dots denote different stages of our TFPI.
In this work, we introduce Thinking-Free Policy Initialization (TFPI), a simple yet effective adaptation to RLVR that bridges long Chain-of-Thought (CoT) distillation and standard RLVR. We first observe that a ThinkingFree operation—explicitly discarding the thinking content via a direct </think> append—can substantially reduce token usage during inference (Section 3.1). Training with ThinkingFree‑adapted inputs improves performance and lowers token consumption, even when evaluated in the original slow‑thinking mode (Section 3.2). We then formally define TFPI in Section 3.3. As illustrated in Figure 1, TFPI accelerates RL convergence, achieves a higher performance ceiling, and produces more token‑efficient reasoning models without requiring specialized rewards or complex training designs (Section 4).
Our contributions are as follows: ❶ We find that ThinkingFree can substantially reduce inference costs for distilled LRMs, and that training with ThinkingFree‑adapted inputs enhances slow‑thinking capability. ❷ We propose TFPI, a fast, low‑cost initialization phase for long‑CoT RL that accelerates RL convergence and generalizes across domains, even when trained solely on mathematics. ❸ We show that long‑CoT RL following TFPI achieves a higher performance ceiling while producing more token‑efficient LRMs without the need for specialized rewards or complex training designs, offering an effective and efficient route to training high‑performing LRMs. ❹ We provide both behavioral and parameter‑level analyses and reveal that TFPI not only preserves the slow‑thinking reasoning pattern but also enables substantially faster rollouts in subsequent long‑CoT RL stages.
2 Preliminary
Notation. In this paper, we define an LLM parameterized by $\theta$ as a policy $\pi_{\theta}$. Let $x$ denote a query and $\mathcal{D}$ the set of queries. Given a response $y$ to a query $x$, its likelihood under the policy $\pi_{\theta}$ is expressed as $\pi_{\theta}(y\mid x)=\prod_{t=1}^{|y|}\pi_{\theta}\left(y_{t}\mid x,y_{<t}\right),$ where $|y|$ denotes the number of tokens in $y$. A query–response pair $(x,y)$ is scored by a rule-based outcome reward $r(x,y)\in\{0,1\}$, indicating whether the response $y$ aligns with the ground truth of $x$.
Proximal Policy Optimization (PPO). PPO (Schulman et al., 2017) constrains the policy update within a proximal region of the old policy $\pi_{\theta_{\text{old}}}$ through the clipping mechanism. Specifically, PPO employs the following objective (we omit the KL regularization term hereinafter for brevity):
$\mathcal{J}_{\text{PPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left[\frac{1}{|y|}\sum_{t=1}^{|y|}\min\left(r_{t}(\theta)\widehat{A}_{t},\,\mathrm{clip}\left(r_{t}(\theta),1-{\varepsilon},1+{\varepsilon}\right)\widehat{A}_{t}\right)\right],$ (1)
with the importance ratio of the token $y_{t}$ is defined as $r_{t}(\theta)=\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|x,y_{<t})}$, the advantage $\widehat{A}_{t}$ of $y_{t}$ is estimated by a value model, and $\varepsilon$ is the clipping range of importance ratios.
Group Relative Policy Optimization (GRPO). GRPO (Shao et al., 2024) bypasses the need for the value model by computing the relative advantage of each response within a group of responses to the same query. Specifically, GRPO optimizes $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{GRPO}}(\theta,x)\right],$ where:
$\mathcal{J}_{\text{GRPO}}(\theta,x)=\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\left(r_{i,t}(\theta)\widehat{A}_{i,t},\,\mathrm{clip}\left(r_{i,t}(\theta),1-{\varepsilon},1+{\varepsilon}\right)\widehat{A}_{i,t}\right)\right].$ (2)
Here $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)$}, $G$ is the number of generated responses to each query $x$ (i.e., the group size), and the importance ratio $r_{i,t}(\theta)$ and advantage $\widehat{A}_{i,t}$ of token $y_{i,t}$ are:
$r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}|x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}|x,y_{i,<t})},\quad\widehat{A}_{i,t}=\widehat{A}_{i}=\frac{r(x,y_{i})-\mathrm{mean}\left(\{r(x,y_{i})\}_{i=1}^{G}\right)}{\mathrm{std}\left(\{r(x,y_{i})\}_{i=1}^{G}\right)},$ (3)
respectively, where all the tokens in $y_{i}$ share the same advantage as $\widehat{A}_{i}$.
Numerous variants have been proposed to improve GRPO, and any RLVR algorithm can be applied to our proposed TFPI stage. In all our experiments, we adopt one widely used variant, DAPO (Yu et al., 2025), as our RLVR algorithm for fair comparison. Details are provided in Appendix A.1.
3 Methodology
3.1 Thinking-Free Mode Enables More Efficient Reasoning
To generate a response $y\sim\pi_{\theta}(\cdot\mid x)$ for a query $x$ using an SFT-distilled LRM $\pi_{\theta}$, the query is typically formatted with a chat template. Template 1 illustrates the template adopted by the Qwen model family (Yang et al., 2025a). We define ThinkingFree as an operator that transforms an input query $x$ into a modified query $x^{\prime}=\textit{ThinkingFree}{}(x)$, in which the thinking content is explicitly omitted. Under this transformation, response generation follows $y\sim\pi_{\theta}(\cdot\mid x^{\prime})$. This mechanism provides explicit control over whether reasoning content is present or absent in the generated output (see Template 2). Additional examples using other chat templates are provided in Appendix A.2. Hereinafter, we use $x$ to denote a query formatted with the thinking template (e.g., Template 1), and $x^{\prime}$ or $\textit{ThinkingFree}(x)$ to denote the corresponding thinking-free version (e.g., Template 2).
Template 1 (Thinking Mode)
[…]
Template 2 (Thinking-Free Mode)
[…]
During inference, converting $x$ into its thinking-free version $x^{\prime}$ can substantially reduce token consumption. To assess how this transformation affects the reasoning capability of SFT-distilled LRMs, we evaluate DeepSeek-Distilled-Qwen-1.5B (abbreviated as DS-1.5B) and Qwen3-4B on the AIME25. Detailed experimental setup is delayed to Appendix B.1. As shown in Figure 2 (Left), applying the ThinkingFree reduces the number of output tokens by more than 70% for both models. It is worth noting that Qwen3-4B is a fast–slow fusion model, whereas DS-1.5B is an SFT-distilled long-CoT model; nevertheless, both exhibit the same trend.
Figure 2: Results of the meta-experiment on the ThinkingFree operation. Left: Average output tokens in thinking mode and ThinkingFree mode on AIME25. Right: Evolution of avg@32 and average output tokens on AIME24 with thinking-mode evaluation over training steps.
3.2 Thinking-Free Training Is Beneficial to Slow-Thinking
To train an SFT-distilled LRM, the training response length should not be too short (Setlur et al., 2025), as this is detrimental to testing performance (An et al., 2025). As evidenced by the dotted line in Figure 2 (Right), training Qwen-3-4B with a 4K response length using DAPO indeed leads to a substantial drop in performance on AIME25. Given that using the ThinkingFree variant for inference can significantly reduce token consumption, we pose an audacious question: can we apply the ThinkingFree operation to all input queries during the rollout stage of RLVR? Moreover, will this approach be beneficial to preserving the original slow-thinking capability of the trained LLM?
Surprisingly, RL trained with ThinkingFree rollout can slightly improve accuracy and reduce token consumption when evaluated in thinking mode, even with very short training context lengths. Figure 2 (Right) illustrates the training dynamics of Qwen-3-4B trained with a 4K response length under the ThinkingFree transformation of queries (detailed settings are in Appendix B.2). Even with a 4K output length, ThinkingFree RL increases the accuracy on AIME25 in thinking mode by approximately 2% , while reducing output tokens by around 20%. In contrast, standard RLVR with a 4K length reduces avg@32 by more than 40%. These results suggest that applying ThinkingFree during rollout can yield steady improvements with minimal training compute.
3.3 Thinking Free Policy Initialization
Hu et al. (2025) introduce a pre-pretraining stage using formal language to accelerate convergence of pretraining, while Wang et al. (2025b) propose a mid-training stage between pretraining and RL-zero to facilitate RLVR. Inspired by these works, we ask: can a dedicated stage for SFT-distilled LRMs improve the efficiency and effectiveness of standard RLVR scaling? ThinkingFree RL, which enhances slow-thinking within short training context windows, requires substantially less compute than standard RLVR. Initializing RLVR with a ThinkingFree RL policy could therefore reduce rollout tokens while achieving stronger downstream performance. We term this step as Thinking Free Policy Initialization (TFPI), a stage preceding standard RLVR for SFT-distilled LRMs that aims to lower rollout costs, raise the ceiling of reasoning ability, and accelerate convergence.
Consider the RLVR objective $\mathcal{J}_{\text{RLVR}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{RLVR}}(\theta,x)\right],$ where $\mathcal{J}_{\text{RLVR}}(\theta,x)$ denotes the per-example objective of any RLVR algorithm (e.g., GRPO in eq. 2 or DAPO in eq. 5). For the TFPI stage, we use a modified objective: $\mathcal{J}_{\text{{TFPI}}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{RLVR}}(\theta,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}})\right],$ where $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}=\textit{ThinkingFree}(x)$}. In the rollout stage of TFPI, $G$ responses are generated conditioned on $x^{\prime}$: $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}}).$ The importance ratios and advantages (eq. 3) are then adapted as:
$r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}},y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}},y_{i,<t})},\quad\widehat{A}_{i,t}=\widehat{A}_{i}=\frac{r({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}},y_{i})-\mathrm{mean}\!\left(\{r({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}},y_{j})\}_{j=1}^{G}\right)}{\mathrm{std}\!\left(\{r({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x^{\prime}},y_{j})\}_{j=1}^{G}\right)},$ (4)
where $r(x^{\prime},y)=r(x,y)$ since the ThinkingFree operator does not alter the ground-truth answer of the original problem. In our experiments, we instantiate RLVR with DAPO, i.e., $\mathcal{J}_{\text{RLVR}}(\theta)=\mathcal{J}_{\text{DAPO}}(\theta)$ (see eq. 5). The details of DAPO (Yu et al., 2025) algorithm is given in Appendix A.1
4 Experiments
4.1 Experimental Settings
In this section, we provide a brief overview of the key experimental setup, including training procedures, baselines, and evaluation details. Additional information can be found in Appendix C.
[…]
Baselines. We compare TFPI with direct RLVR training from an SFT-distilled LRM (“Direct RL”) under matched total training compute (Table 1). To assess TFPI as a pre-RLVR stage, we run “TFPI + RL” with similar compute and compare against “Direct RL” (Table 2), also including competitive LRMs of the same size, such as Polaris (An et al., 2025), DeepScaleR (Luo et al., 2025b), Skywork-OR1 (He et al., 2025), and so on. For efficiency analysis, we compare with several RL-based efficient reasoning baselines (Table 3) under the same evaluation for fair comparison, including TLMRE (Arora & Zanette, 2025), AdaptThink (Zhang et al., 2025b), AutoThink (Tu et al., 2025), Laser (Liu et al., 2025a), L1Max (Aggarwal & Welleck, 2025), and ThinkLess (Fang et al., 2025).
Evaluation Details. Our evaluation benchmarks cover: ❶ Math Reasoning: AIME24/25 and BeyondAIME (ByteDance-Seed, 2025). ❷ Multi-Task Reasoning: GPQA-Diamond (Rein et al., 2024). ❸ Code Generation: LiveCodeBench (Jain et al., 2024). ❹ Instruction Following: IFEval (Zhou et al., 2023). ❷ ❸ ❹ are out-of-domain evaluation. Following Guo et al. (2025), we generate multiple outputs (ranging from 4 to 32 depending on the size of the test set) per problem and report pass@1 accuracy. Note that ❶ is in-domain evaluation, and For IFEval, we report the strict prompt accuracy. All evaluation scripts are adapted from the DeepscaleR codebase (Luo et al., 2025b), with detailed decoding parameters provided in Appendix C.3.
4.2 TFPI Enhances the Slow-Thinking of Distilled Reasoning Models
Table 1: Results of TFPI vs. direct RL across different benchmarks. “Avg@k” denotes the average accuracy (%) over $k$ random generations (i.e., pass@1). All models are evaluated in thinking mode. The total training compute for the 3 stages of TFPI equals that of “direct RL” for fair comparison. Darker colors in the cell background denote better results within each model group.
Models Mathematics Multi-Task Code Instruction Overall
AIME 24 AIME 25 Beyond AIME GPQA LiveCode IFEval Overall
Avg@32 Avg@32 Avg@8 Avg@8 Avg@8 Avg@4 Avg.
DeepSeek-Distill-Qwen-1.5B
Initial Model 29.6 23.0 8.7 16.3 17.7 36.6 22.0
- Direct RL 33.9 27.1 11.9 19.2 18.2 41.8 25.3
- TFPI stage 1 32.1 26.9 13.9 29.3 18.4 39.5 26.7
- TFPI stage 2 36.8 28.4 14.5 27.8 20.5 42.3 28.4
- TFPI stage 3 40.1 30.8 13.8 29.6 19.9 40.8 29.2
Qwen3-4B
Initial Model 73.6 68.3 43.4 56.8 54.9 64.9 60.3
Direct RL 75.7 67.0 43.6 56.3 52.5 66.0 60.2
TFPI stage 1 75.2 67.8 42.4 57.9 55.3 66.0 60.8
TFPI stage 2 76.0 68.2 44.7 57.8 54.8 64.8 61.0
TFPI stage 3 79.9 70.6 46.7 58.5 57.0 70.2 63.8
DeepSeek-Distill-Qwen-7B
Initial Model 56.3 40.0 25.0 36.9 39.5 55.3 42.2
- Direct RL 57.9 40.4 26.6 38.0 38.3 56.7 43.0
- TFPI stage 1 59.4 40.4 29.2 49.0 39.6 56.2 45.6
- TFPI stage 2 61.8 43.9 31.5 48.0 42.0 57.1 47.4
- TFPI stage 3 62.0 44.6 31.1 46.8 42.1 60.2 47.8
To evaluate the impact of TFPI on the slow-thinking capabilities, we present the results of TFPI versus “Direct RL” under the same training compute in Table 1. From the table, we have:
TFPI substantially enhances the slow-thinking capabilities of distilled LRMs even when trained with a small response length. For example, on DS-1.5B, TFPI Stage 1 raises the overall average accuracy from $22.0\%$ to $26.7\%$ (+4.7%) despite being restricted to a 2K training response length. Results of DS-1.5B continue to improve through stages 1 to 3 on AIME25, accuracy increases from 23.023.0% (initial model) to 26.926.9% after Stage 1, 28.428.4% after Stage 2, and 30.830.8% after Stage 3, representing a total gain of +7.8%. Similar improvements are observed for Qwen3-4B and DS-7B. These findings indicate that TFPI enables effective training under low-cost settings (short context windows), and that combining this with multi-stage RL yields substantial accuracy gains.
Compared with “Direct RL”, TFPI delivers faster and larger accuracy improvements under the same training cost. TFPI outperforms “Direct RL” in nearly all configurations. For example, Qwen3-4B with TFPI attains $63.8\%$ overall accuracy versus $60.2\%$ for “Direct RL” (+3.6%), while DS-7B improves by +4.8% ($47.8\%$ vs. $43.0\%$). Given that the equal training compute, these results imply that TFPI achieves convergence more efficiently than conventional long-CoT RL training.
Improvements of TFPI exhibit some degree of generalizability across domains, even when trained exclusively on mathematics. Although TFPI is trained solely on Polaris-53K (math-specific data), it demonstrates great out-of-domain improvements as well. For example, on DeepSeek-Distill-Qwen-1.5B, GPQA accuracy increases from $16.3\%$ to $29.6\%$, LiveCodeBench from $17.7\%$ to $19.9\%$, and IFEval from $36.6\%$ to $40.8\%$ after Stage 3. Notably, improvements on mathematical benchmarks are often consistent across successive training stages, whereas other domains sometimes exhibit fluctuations (e.g., GPQA for DS-7B and LiveCodeBench for DS-1.5B). This suggests that incorporating more diverse training data spanning multiple domains could be helpful for TFPI.
4.3 TFPI as a Foundation for RLVR to Achieve Higher Performance
Table 2: Results (%) of RL after TFPI (“TFPI+RL”) vs. “Direct RL” across different benchmarks. “Avg@k” denotes the average accuracy (%) over $k$ random generations (i.e., pass@1). For LRMs marked with “*”, results are taken from the corresponding reports (see Appendix C.4); results of 4B models are from our own runs with 48K response length. All models are evaluated in thinking mode. The total training compute for “TFPI+RL” is matched to that of “Direct RL” for fair comparison. Darker colors in the cell background denote better results.
Models Mathematics Multi-Task Code Instruction Overall
AIME 24 AIME 25 Beyond AIME GPQA LiveCode IFEval Overall
Avg@32 Avg@32 Avg@8 Avg@8 Avg@8 Avg@4 Avg.
DeepSeek R1 79.8 65.0 42.4 71.5 64.3 86.1 68.2
Seed-1.5-Thinking 86.7 74.0 48.0 77.3 64.9 87.4 73.0
Claude4 Opus Thinking - 75.5 - 79.6 48.9 89.7 -
Qwen3-235B-Thinking 85.7 81.5 - 71.1 55.7 83.4 -
Qwen3-4B Direct RL 78.8 71.5 46.4 56.2 54.3 65.1 62.0
Qwen3-4B TFPI stage 3 79.9 70.6 46.7 58.5 57.0 70.2 63.8
Qwen3-4B TFPI + RL 80.8 76.0 49.5 61.1 55.7 71.3 65.7
Qwen3-4B-2507-Thinking 84.7 79.2 51.5 66.4 62.4 68.0 68.7
- TFPI only 89.0 81.2 51.3 70.1 65.5 66.7 70.6
Table 2 shows the representative results between “TFPI + RL” with “Direct RL” under the same training cost (full results in Table 7 in Appendix E). The results lead to the following observations:
TFPI can raise the upper bound of RL-trained performance. RL training on a TFPI-trained model can still yield notable improvements, particularly in mathematics. For example, adding an RL stage after TFPI for Qwen3-4B boosts AIME25 accuracy from $70.6\%$ to $76.0\%$ ($5.4\%$). Similarly, Beyond AIME scores increase by more than $2\%$. Across different model scales, applying RL after TFPI consistently achieves higher accuracies than “Direct RL” under the same compute budget. For instance, on Qwen3-4B, the overall accuracy rises from $62.0\%$ (Direct RL) to $65.7\%$ (TFPI+RL).These results suggest that incorporating TFPI as an intermediate stage between long CoT distillation and standard RLVR can be beneficial for elevating final performance.
[…]
4.4 TFPI Improves the Token Efficiency of Distilled Reasoning Models
Table 3: Comparison of the Thinking-Free inference mode of TFPI with efficient reasoning baselines across various reasoning tasks. “Avg@k” denotes the average accuracy (in %) over k generations (i.e., pass@1), and “Toks” indicates the average output length in thousands of tokens (K). Models with “*” are trained from DeepScaleR-1.5B, while the remaining are from DS-1.5B. Darker colors in the cell background denote better results.
Models AIME 24 AIME 25 Beyond AIME GPQA Overall
avg@32 Toks avg@32 Toks avg@8 Toks avg@8 Toks Avg. Toks
1) TLMRE-DS-1.5B ($\alpha=0.1$) 27.6 12.9 24.8 12.5 9.2 11.9 14.8 7.5 19.1 11.2
2) AdaptThink-1.5B ($\delta=0.05$) 28.1 8.0 22.6 7.9 10.0 4.8 14.8 5.1 18.9 6.5
3) AutoThink-DS-1.5B-Stage3 30.3 10.2 25.2 9.1 9.4 8.7 17.4 7.0 20.6 8.7
4) Laser-DE-L4096-1.5B 30.3 8.3 24.9 7.4 9.7 7.4 21.1 4.7 21.5 6.9
5) AutoThink-Stage3 38.9 8.7 28.9 7.7 11.6 7.8 27.3 5.4 26.7 7.4
6) L1-1.5B-Max 27.2 3.2 26.3 2.9 9.1 3.1 32.4 2.3 23.8 2.9
7) Thinkless-1.5B-RL 28.4 11.3 24.1 11.1 8.1 11.7 20.3 12.7 20.2 11.7
DS-1.5B (Thinking) 29.6 16.7 23.0 16.5 8.7 14.4 16.3 9.8 19.4 14.3
DS-1.5B (Thinking-Free) 12.4 5.7 10.9 4.4 4.4 3.4 4.2 0.9 8.0 3.6
- TFPI stage 1 21.9 1.6 15.3 1.4 8.7 1.3 32.9 0.8 19.7 1.3
- TFPI stage 2 31.5 3.4 24.2 3.1 10.1 2.9 35.3 1.6 25.3 2.7
- TFPI stage 3 37.5 5.3 28.4 5.0 12.4 4.9 35.6 2.6 28.5 4.4
We compare the thinking-free inference with other RL-based efficient reasoning baselines in 1.5B size in Table 3 (see also Table 8 in Appendix E). We draw the following conclusions:
Both accuracy and token efficiency steadily improve after stage 3 of TFPI. For DS-1.5B, thinking-free accuracy on AIME24 increases from $29.6\%$ (initial model) to $37.5\%$ (TFPI stage 3), while the average output length remains substantially shorter than that of the original thinking mode (5.3K vs. 16.7K tokens). A similar trend is observed for Qwen3-4B, where accuracy improves from $26.9\%$ to $75.1\%$ across stages, with output lengths still far below those of the original thinking model. These results demonstrate that TFPI naturally produces more token-efficient LRMs, offering an alternative pathway to train models that deliver both high accuracy and reduced output length.
Compared with other RL-based token control methods, TFPI achieves the best performance–efficiency trade-off without specialized reward or training designs. In DS-1.5B, both TFPI stage 2 and stage 3 outperform almost all baselines in terms of overall accuracy while maintaining competitive or lower token usage. We visualize the overall accuracy–token usage trade-off in Figure 1 (Right), where TFPI consistently lies on the Pareto frontier across different stages. This observation motivates a rethinking of existing “token-efficient reasoning RL designs” that rely heavily on specialized length reward shaping: training with TFPI offers an alternative paradigm in which a strong slow-thinking LRM can be obtained, and a more efficient variant can be realized simply by switching to the thinking-free mode without any additional length-control mechanisms.
5 Analysis
5.1 Why TFPI Leads to Better Thinking-Mode Inference?
In this section, we delve deeper into why TFPI, which leverages void thinking content in Template 2, can generalize to enhance reasoning in Template 1. We analyze DS-1.5B from two perspectives:
Behavioural level: The learned verification behaviours after </think> during TFPI can generalize to the slow-thinking verification occurring within the <think> and </think>. The blue lines in Figure 3 show the ratio of verification steps (i.e., the number of verification steps divided by the total number of steps; see Appendix D for details) for the training set (thinking-free mode) and AIME 25 dataset (thinking mode). We observe that the verification ratio exhibits a similar trend: a rapid drop in Stage 1, followed by steady growth in Stage 2, and a sharp increase in Stage 3. Notably, the sharp decline in Stage 1 resembles an information compression process. In Stages 2 and 3, the model begins to explore more extensively (Figure 3 Right), which may explain why TFPI achieves superior performance. As verification is believed to be vital for slow-thinking reasoning (Setlur et al., 2025), the observed generalization of verification behavior suggests a transfer from thinking-free training to inference in thinking mode.
Parameter level: TFPI explores the parameter space more extensively at a faster pace, with its parameter update directions progressively aligning with those of “Direct RL.” As shown in Figure 4 (Left), the PCA visualization of the initial model, TFPI-trained checkpoints, and Direct RL checkpoints exibits distinct trajectories in parameter space. The TFPI begins from the initial model (A), moves through intermediate points (B1), (B2), and (B3), and ultimately converges to a region near the Direct RL final checkpoint (C). This indicates that TFPI traverses a larger and more diverse region of parameter space before reaching a point close to the RL-trained model. Such broad exploration may help explain why TFPI can lead to better LRMs. Furthermore, Figure 4 (Right) shows that the cosine similarity between the TFPI-trained checkpoints and the Direct RL final checkpoint steadily increases across nearly all layers throughout training. This suggests that during TFPI training, the parameter updates share similarities with those in standard long-CoT training.
Figure 3: Behaviour-Level Analysis of DS-1.5B over the TFPI Training Course. The ratio of verification steps and the average output tokens over training steps on the training set in thinking-free mode (Left) and on AIME25 in thinking mode (Right) in 3 stages of TFPI.
Figure 4: Parameter-Level Analysis. Left: PCA projection of model parameters from DS-1.5B to final checkpoints. TFPI (blue) starts at A, passes through intermediate points (B1, B2, B3), and ultimately converges near the Direct RL final checkpoint (C). Right: Cosine similarity between parameter updates of TFPI-trained checkpoints and (C-A) across layers during training.
5.2 About Reasoning Pattern and Rollout Speed
TFPI Preserves the Reasoning Pattern of Thinking Mode. In thinking mode (Template 1), a response $y$ comprises a long thinking section and a concise answer $y^{\text{ans}}$. In ThinkingFree mode (Template 2), the thinking section is omitted, and $y^{\text{ans}}$ contains a shorter reasoning path. While standard RL tends to lengthen the thinking part, TFPI increases the length of $y^{\text{ans}}$ due to the absence of explicit thinking. As shown in Figure 5 (Left), for DS-1.5B trained with TFPI and evaluated in thinking mode, $|y^{\text{ans}}|$ remains stable at 500500580580 tokens, whereas $|y^{\text{ans}}|/|y|$ rises as the total length $|y|$ decreases. This indicates that TFPI preserves the core reasoning pattern of slow-thinking rather than drifting toward an excessively extended “slow-slow thinking” behavior.
TFPI Speeds Up Rollout for Long-CoT RL Training. Another advantage of TFPI is its ability to speed up the rollout stage in standard long-CoT RL training. As shown in Figure 5 (Right), when directly performing RL from DS-1.5B, the average output tokens during rollout on the training set start at over 9K, decrease to around 7.5K within the first 300 steps, and then fluctuate around 7.5K in the later stages. In contrast, when RL is performed after the TFPI phase, the average output tokens start at only 6K and the maximum length is below 7K tokens.
Figure 5: Left: For TFPI with DS-1.5B on AIME25 in thinking mode, showing the average number of answer tokens (excluding the thinking part) and the ratio of answer length to total response length over training steps. Right: For long CoT RL with DS-1.5B, showing the average number of tokens during rollout on the training set over training steps, with and without the TFPI stage.
6 Related Work
LRMs & Efficiency RLVR. RLVR has enabled the development of numerous high-performing large reasoning models (LRMs) (Yang et al., 2025a; Zeng et al., 2025a; An et al., 2025), inspiring research on model behaviors (Liu et al., 2025b; Wang et al., 2025a), novel algorithms (Liu et al., 2025b; Zheng et al., 2025), multimodal extensions (Meng et al., 2025; Xiao et al., 2025; Wang et al., 2025c), tool integration (Jin et al., 2025; Feng et al., 2025; Li et al., 2025a; Xue et al., 2025; Team et al., 2025a), and other related directions (Xu et al., 2025a; Zhang et al., 2025a; Chen et al., 2025c; Zhang et al., 2025c). RLVR can be applied directly to base LLMs (“RL zero”) (Zeng et al., 2025b; Guo et al., 2025) or initialized from SFT-distilled long-CoT models, the latter typically yielding stronger results (Luo et al., 2025b; An et al., 2025). A major challenge for RLVR is the cost of training with long contexts, as longer outputs are often necessary for harder tasks (Shrivastava et al., 2025; Zeng et al., 2025a), consuming high computational resources. Multi-stage RLVR mitigates this by starting with shorter contexts and gradually extending them (Luo et al., 2025b; An et al., 2025; He et al., 2025), while algorithmic approaches modify GRPO to reduce length bias (Yu et al., 2025; Liu et al., 2025b; Wang et al., 2025a). Orthogonal to these strategies, we introduce TFPI as a lightweight stage before RLVR, improving efficiency and strengthening the slow-thinking mode at inference with minimal training cost, thus facilitating more effective subsequent RLVR.
Efficient Reasoning. To address the issue of overthinking (Chen et al., 2024; Sui et al., 2025), considerable efforts have been made, including prompt-based methods (Muennighoff et al., 2025; Yang et al., 2025c; Fu et al., 2025a; Chen et al., 2025a; Fu et al., 2025b), SFT-based approaches (Kang et al., 2025; Ma et al., 2025; Munkhbat et al., 2025; Luo et al., 2025a), and RL-related designs. RL-related approaches can be further categorized into length-based reward shaping (Team et al., 2025b; Aggarwal & Welleck, 2025; Arora & Zanette, 2025; Liu et al., 2025a), integration of fast and slow thinking (Fang et al., 2025; Zhang et al., 2025b; Lou et al., 2025; Tu et al., 2025; Jiang et al., 2025; Zhang et al., 2025d), and thinking budget control (Li et al., 2025b; Hammoud et al., 2025; Wen et al., 2025). These methods primarily trade accuracy for efficiency and rely on specialized reward functions or training strategies to encourage more efficient reasoning. In contrast, our proposed TFPI naturally yields even more efficient LRMs without specialized rewards or training designs.
7 Conclusion
After recognizing the benefits of ThinkingFree for both inference and the training of distilled reasoning models, we introduce TFPI, a cost‑efficient intermediate stage between long‑CoT distillation and standard RL training. As a strong initialization point, TFPI accelerates RL convergence, enhances attainable performance, and promotes more token‑efficient reasoning without complex reward shaping or elaborate training pipelines. We further explain the factors behind TFPI’s success from both the behavioral and parameter levels. Overall, TFPI provides a complementary path for building “token‑efficient” LRMs, offering an effective and efficient alternative to current RL paradigms.
Appendix A Background and Preliminary
A.1 RLVR Algorithms
Numerous variants have been proposed to improve GRPO. For example, DAPO (Yu et al., 2025) introduces token-level normalization and dynamic sampling; Dr GRPO (Liu et al., 2025b) removes length bias to prevent incorrect responses from growing longer over time; and Wang et al. (2025a) train selectively on forking tokens (see also Section 6). Our TFPI is orthogonal to these RLVR algorithms. That is to say, any RLVR algorithm can be applied to our proposed TFPI stage. To mitigate the effect of RLVR algorithms, we employ DAPO as our RLVR algorithm in all experiments for fair comparison.
Dynamic sAmpling Policy Optimization (DAPO)
Building on GRPO, DAPO (Yu et al., 2025) introduces a clip-higher mechanism, incorporates dynamic sampling and applies a token-level policy gradient loss. The objective is given by $\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{DAPO}}(\theta,x)\right],$ where:
$\mathcal{J}_{\text{DAPO}}(\theta,x)=$ $\Bigg[\frac{1}{\sum_{i=1}^{G}|y_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|y_{i}|}\min\Big(r_{i,t}(\theta)\,\widehat{A}_{i,t},\,\mathrm{clip}\big(r_{i,t}(\theta),1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}}\big)\,\widehat{A}_{i,t}\Big)\Bigg],$
$\text{s.t.}\quad 0<\left|\left\{y_{i}\mid\texttt{is\_equivalent}(y_{i},x)\right\}\right|<G,$ (5)
where $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)$} and $G$ is the number of generated responses to each query $x$ (i.e., the group size) and $r_{i,t}(\theta),\,\widehat{A}_{i,t}$ is computed as in eq. 3.
A.2 ThinkingFree Operation
In Section 3.1, we have shown chat templates for both the original query $x$ and its thinking-free version $x^{\prime}=\textit{ThinkingFree}(x)$ for Qwen models. Here, we showcase one additional example under the DeepSeek (Guo et al., 2025) template as below.
Template 3 (Thinking Mode (DeepSeek))
[…]
Template 4 (Thinking-Free Mode (DeepSeek))
[…]
Appendix B Meta Experiments
B.1 Token Consumption of ThinkingFree
We provide the detailed experimental setup for the meta-experiment in Section 3.1 in this appendix.
For DS-1.5B, we use the decoding parameters suggested by Guo et al. (2025) in thinking mode: temperature = 0.6, top-$p$ = 0.95, and top-$k$ = 1-1. For Qwen3-4B, we adopt the recommended settings from Yang et al. (2025a) in thinking mode: temperature = 0.6, top-$p$ = 0.95, and top-$k$ = $20$. For ThinkingFree, we set temperature = 0.7, top-$p$ = 0.8, and use both top-$k$ = 1-1 and top-$k$ = $20$. The maximum output length is fixed at 32K tokens for both modes. For evaluation, we sample 32 generations per query on AIME 2025 and report the average number of output tokens. The parameters are given in Table 4 and the results are shown in Figure 2 (Left).
Table 4: Decoding Parameters of Meta-Experiment in Section 3.1
Thinking Mode Thinking-Free Mode
DS-1.5B $T=0.6$, top-$p$ = 0.95, and top-$k$ = 1-1 $T=0.7$, top-$p$ = 0.8, and top-$k$ = 2020
Qwen3-4B $T=0.6$, top-$p$ = 0.95, and top-$k$ = 2020 $T=0.7$, top-$p$ = 0.8, and top-$k$ = 2020
B.2 Detailed Setup of ThinkingFree Training
We provide the detailed experimental setup for the meta-experiment discussed in Section 3.2 in this appendix.
For training, we use the DAPO (Yu et al., 2025) algorithm, with configurations identical to those in Appendix C.1, except that the maximum output length is set to 4K. For evaluation, we set the maximum output length to 48K and perform testing in thinking mode as described in Template 1. All other evaluation parameters follow Appendix C.3. Note that the initial avg@32 value in Figure 2 (Right) is higher than the value reported in (Yang et al., 2025a) (68.2 vs. 65.6), because we adopt the RoPE scaling method described in Polaris (An et al., 2025).
Appendix C Experimental Details
In this appendix, we provide the details of our main experiments in Section 4.
C.1 Training Details
We build on the VeRL codebase (Sheng et al., 2024), with the RLVR loss following the DAPO recipe (Yu et al., 2025). Specifically, the RLVR loss is defined in eq. 5. For our TFPI, it becomes
$\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{DAPO}}(\theta,x^{\prime})\right],\quad x^{\prime}=\textit{ThinkingFree}(x).$ (6)
For fair comparison, we use identical hyperparameters across methods. For clip-higher, we set $\varepsilon_{\text{low}}=0.2$ and $\varepsilon_{\text{high}}=0.28$. Training is performed with a batch size and mini-batch size of 256, a learning rate of $10^{-6}$, and no warm-up scheduling. Both KL divergence loss and entropy loss are excluded.
For rollout, we use temperature =1=1, $\text{topp}=1$, and $\text{topk}=-1$. We generate 8 rollouts per problem. Experiments are conducted on DS-1.5B, Qwen3-4B, and DS-7B, with Polaris-53K (An et al., 2025) as the training dataset. In principle, we could apply dataset filtering at each training stage to accelerate training An et al. (2025). However, for fairness, we deliberately use the full training set for all experiments.
For Direct RLVR, the maximum output length is set to 16K for DS-1.5B and DS-7B, and 32K for Qwen3-4B. For TFPI, we adopt a multi-stage training strategy (An et al., 2025; Luo et al., 2025b):
  • DS-1.5B and DS-7B: $2\text{K}\rightarrow 4\text{K}\rightarrow 8\text{K}$.
  • Qwen3-4B: $4\text{K}\rightarrow 8\text{K}\rightarrow 16\text{K}$.
Our experiments are conducted with 32 H20 GPUs. A summary of the number of training steps and training time are provided in Table 5.
Table 5: Training Steps and Time of Main Experiments. “kh” denotes one thousand H20 Hours.
DS-1.5B Qwen3-4B DS-7B DS-1.5B Qwen3-4B DS-7B
TFPI Stage 1 2K, 1K steps 4K, 100 steps 2K, 1K steps 0.84 kh 0.35 kh 1.73 kh
TFPI Stage 2 4K, 440 steps 8K, 56 steps 4K, 232 steps 0.62 kh 0.52 kh 0.82 kh
TFPI Stage 3 8K, 440 steps 16K, 64 steps 8K, 144 steps 1.21 kh 0.91 kh 0.94 kh
TFPI Total - - - 2.66 kh 1.79 kh 3.50 kh
Direct RLVR 16K, 456 steps 32K, 20 steps 16K, 280 steps 2.67 kh 1.9 kh 3.52 kh
TFPI: RLVR 16K, 472 steps 32K, 192 steps 536 steps 2.49 kh 13.5 kh 6.55 kh
TFPI + RLVR Total - - - 5.16 kh 15.4 kh 10 kh
Direct RLVR Total 16K, 896 steps 32K, 216 steps 16K, 820 steps 5.17 kh 15.7 kh 10.1 kh
C.2 Baselines
To evaluate the efficacy of TFPI, we compare TFPI with direct RLVR training from an SFT-distilled LRM (“Direct RL” for short) under the same total training compute, i.e., the combined compute of the three TFPI stages equals that of direct RLVR (Table 1). To examine the effect of inserting a TFPI stage before RLVR, we apply TFPI to an SFT-distilled model (“TFPI + RL”), continue with standard RLVR, and compare the results with “Direct RL” using approximately the same training compute (Table 2). We also include several high-performing LRMs of the same model size from previous works for reference, including Polaris (An et al., 2025), DeepScaleR (Luo et al., 2025b), Skywork-OR1 (He et al., 2025), and AReal-RL. For both Table 1 and Table 2, all models are evaluated in thinking mode (see also Appendix C.3). To assess the impact of TFPI on reasoning efficiency, we compare our TFPI-trained model with various RL-based efficient reasoning baselines, including TLMRE (Arora & Zanette, 2025), AdaptThink (Zhang et al., 2025b), AutoThink (Tu et al., 2025), Laser (Liu et al., 2025a), L1Max (Aggarwal & Welleck, 2025), and ThinkLess (Fang et al., 2025) (Table 3). The training and testing settings of these baselines, as reported in their original papers, are summarized in Table 6. For fair comparison, we standardize the testing parameters to topp = 0.95, topk = 1-1, and $T$ = 0.6 with a 32K maximum length, following the recommendations of DeepSeek-R1 (Guo et al., 2025).
Table 6: Training and evaluation details of efficient reasoning baselines from original papers. We unify the evaluation setting for fair comparison in Table 3. Details are provided in Appendix C.2 and C.3.
Model Training Data Training Length Test Length Evaluation Details
DeepSeek-Distill-Qwen-1.5B
TLMRE-DS-1.5B ($\alpha=0.1$) 3.2K examples from NuminaMath 32K 10 generations for AIME24
AdaptThink-1.5B ($\delta=0.05$) DeepScaleR-40K 16K 16K $T=0.6$, 16 generations for AIME24
AutoThink-DS-1.5B-Stage3 DeepScaleR-40K 8K $\rightarrow$ 16K $\rightarrow$ 24K $T=0.6$, 16 generations
Laser-DE-L4096-1.5B DeepScaleR-40K 32K 16 generations for AIME24
DeepscaleR-1.5B-Preview
AutoThink-Stage3 DeepScaleR-40K 8K $\rightarrow$ 16K $\rightarrow$ 24K $T=0.6$, 16 generations
L1-1.5B-Max DeepScaleR-40K 4K 8K
Thinkless-1.5B-RL DeepScaleR-40K 24K
C.3 Evaluation Details
To comprehensively evaluate model capabilities, we employ a diverse set of benchmarks covering mathematical reasoning, multi-task reasoning, code generation, and instruction-following:
  1. 1.
    Mathematical reasoning: We evaluate on AIME24, AIME25, and BeyondAIME (ByteDance-Seed, 2025). For AIME24 and AIME25, we report pass@1 accuracy using 32 samples per problem (avg@32); for BeyondAIME, we report avg@8.
  2. 2.
    Multi-task reasoning: We evaluate on GPQA-Diamond (Rein et al., 2024) and report pass@1 with 8 samples per problem.
  3. 3.
    Code generation: We assess coding ability on LiveCodeBench (Jain et al., 2024) (2024-08–2025-01 subset, aligned with DeepSeek-R1 (Guo et al., 2025)), reporting pass@1 with 8 samples per problem.
  4. 4.
    Instruction-following: We evaluate on IFEval (Zhou et al., 2023) and report pass@1 of the strict prompt accuracy with 4 samples per problem.
All evaluation codes are adapted from the DeepscaleR (Luo et al., 2025b) codebase, where vLLM (Kwon et al., 2023) is leveraged to accelerate inference. For IFEval, we use the same codes provided by the official paper (Zhou et al., 2023).
We provide our decoding parameters as follows:
  • Table 1: We set the temperature to 0.6 and topp = 0.95. For LRMs trained from DS-1.5B and DS-7B, we use topk = 1-1 with a maximum sequence length of 32K tokens. For LRMs trained from Qwen3-4B, we use topk = $20$ with a maximum sequence length of 48K tokens, applying RoPE scaling as proposed by An et al. (2025).
  • Table 2: We set the temperature to 0.6 and topp = 0.95. For LRMs initialized from DS-1.5B and DS-7B, we use topk = 1-1 with a maximum sequence length of 32K tokens. For LRMs initialized from Qwen3-4B, we use topk = $20$ with a maximum sequence length of 48K tokens, again applying RoPE scaling as proposed by An et al. (2025). For models marked with “*”, we report the results from their original publications (see Appendix C.4).
  • Table 3: For efficient reasoning baselines listed in Table 6 and the thinking mode of initial models, we set topp = 0.95, topk = 1-1, and $T$ = 0.6, with a maximum sequence length of 32K tokens (48K for Qwen3-4B). For the thinking-free mode of initial models and our TFPI, we set topp = 0.8, topk = $20$, and $T$ = 0.7 with a maximum sequence length of 32K tokens, following Yang et al. (2025a).
C.4 Source of Some Results in Table 2
Results for LRMs marked with “*” are taken directly from the Seed-1.5-Thinking report (Seed et al., 2025) and the corresponding Hugging Face page of Qwen3-2507. Note that the LiveCodeBench test set subsets of these results, and the metric of IFEval may differ from those in our experiments; their results are included for reference only.
Appendix D Analysis Details
[…]
Appendix E More Results
Due to page limit, we present only representative results in Sections 4.3 and 4.4, with the complete results provided in Tables 7 and 8, respectively.
Table 7: Results (%) of RL after TFPI (“TFPI+RL”) vs. “Direct RL” across different benchmarks. “Avg@k” denotes the average accuracy (%) over $k$ random generations (i.e., pass@1). For LRMs marked with “*”, results are taken from the corresponding reports (see Appendix C.4); all other results are from our own runs. All models are evaluated in thinking mode. The total training compute for “TFPI+RL” is matched to that of “Direct RL” for fair comparison.
Models Mathematics Multi-Task Code Instruction Overall
AIME 24 AIME 25 Beyond AIME GPQA LiveCode IFEval Overall
Avg@32 Avg@32 Avg@8 Avg@8 Avg@8 Avg@4 Avg.
DeepSeek R1 79.8 65.0 42.4 71.5 64.3 86.1 68.2
Seed-1.5-Thinking 86.7 74.0 48.0 77.3 64.9 87.4 73.0
Claude4 Opus Thinking - 75.5 - 79.6 48.9 89.7 -
Qwen3-235B-Thinking 85.7 81.5 - 71.1 55.7 83.4 -
Qwen3-8B 75.3 67.0 43.2 61.7 51.7 66.0 61.7
DS-R1-0528-Qwen3-8B 82.8 74.6 50.6 61.4 59.0 73.9 67.0
1.5B Size
DeepScaleR-1.5B 37.8 31.6 13.1 19.1 21.9 40.5 27.3
DS-1.5B Direct RL 37.2 28.5 12.9 24.6 19.5 39.5 27.0
DS-1.5B TFPI stage 3 40.1 30.8 13.8 29.6 19.9 40.8 29.2
DS-1.5B TFPI + RL 42.3 32.9 15.1 27.8 20.9 41.0 30.0
4B Size
Polaris-4B-Preview 73.2 70.7 39.9 54.9 39.7 63.9 57.0
Qwen3-4B Direct RL 78.8 71.5 46.4 56.2 54.3 65.1 62.0
Qwen3-4B TFPI stage 3 79.9 70.6 46.7 58.5 57.0 70.2 63.8
Qwen3-4B TFPI + RL 80.8 76.0 49.5 61.1 55.7 71.3 65.7
Qwen3-4B-2507-Thinking 84.7 79.2 51.5 66.4 62.4 68.0 68.7
- TFPI only 89.0 81.2 52.5 70.1 65.5 66.7 70.8
7B Size
AReal-boba-RL-7B 61.5 46.1 30.7 35.4 34.2 57.5 44.2
Skywork-OR1-7B-Preview 61.2 46.4 31.2 35.2 43.3 55.4 45.5
Polaris-7B-Preview 70.6 48.8 37.0 34.1 43.9 55.6 48.3
DS-7B Direct RL 62.3 47.9 30.1 36.9 42.8 57.1 46.2
DS-7B TFPI stage 3 62.0 44.6 31.1 46.8 42.1 60.2 47.8
DS-7B TFPI + RL 65.3 47.1 33.2 45.9 43.3 57.3 48.7
Table 8: Comparison of the Thinking-Free inference mode of TFPI with efficient reasoning baselines across various reasoning tasks. “Avg@k” denotes the average accuracy (in %) over k generations (i.e., pass@1), and “Toks” indicates the average output length in thousands of tokens (K).
Models AIME 24 AIME 25 Beyond AIME GPQA Overall
avg@32 Toks avg@32 Toks avg@8 Toks avg@8 Toks Avg. Toks
DeepSeek-Distill-Qwen-1.5B
TLMRE-DS-1.5B ($\alpha=0.1$) 27.6 12.9 24.8 12.5 9.2 11.9 14.8 7.5 19.1 11.2
AdaptThink-1.5B ($\delta=0.05$) 28.1 8.0 22.6 7.9 10.0 4.8 14.8 5.1 18.9 6.5
AutoThink-DS-1.5B-Stage3 30.3 10.2 25.2 9.1 9.4 8.7 17.4 7.0 20.6 8.7
Laser-DE-L4096-1.5B 30.3 8.3 24.9 7.4 9.7 7.4 21.1 4.7 21.5 6.9
Initial Model (Thinking) 29.6 16.7 23.0 16.5 8.7 14.4 16.3 9.8 19.4 14.3
Initial Model (Thinking-Free) 12.4 5.7 10.9 4.4 4.4 3.4 4.2 0.9 8.0 3.6
- TFPI stage 1 21.9 1.6 15.3 1.4 8.7 1.3 32.9 0.8 19.7 1.3
- TFPI stage 2 31.5 3.4 24.2 3.1 10.1 2.9 35.3 1.6 25.3 2.7
- TFPI stage 3 37.5 5.3 28.4 5.0 12.4 4.9 35.6 2.6 28.5 4.4
DeepscaleR-1.5B-Preview
AutoThink-Stage3 38.9 8.7 28.9 7.7 11.6 7.8 27.3 5.4 26.7 7.4
L1-1.5B-Max 27.2 3.2 26.3 2.9 9.1 3.1 32.4 2.3 23.8 2.9
Thinkless-1.5B-RL 28.4 11.3 24.1 11.1 8.1 11.7 20.3 12.7 20.2 11.7
Qwen3-4B
Initial Model (Thinking) 73.6 18.0 68.3 22.3 43.4 23.5 56.8 10.7 60.5 18.6
Initial Model (Thinking-Free) 26.9 7.2 20.2 5.5 11.0 4.2 45.0 2.3 25.8 4.8
- TFPI stage 1 43.9 3.7 31.1 3.4 23.2 3.4 47.8 1.5 36.5 3.0
- TFPI stage 2 63.7 7.2 52.8 8.0 33.0 7.0 50.9 2.3 50.1 6.1
- TFPI stage 3 75.1 10.5 66.9 12.7 42.1 12.7 55.9 3.7 60.0 9.9
Qwen3-4B-Instruct-2507 60.4 7.9 46.1 7.4 33.0 7.3 64.4 4.8 51.0 6.9